Write the dynamic model in eqn. 1 in terms of the deviation variables 0*(t), q*(t), and 0*(t)


1
Expert's answer
2022-02-19T06:53:02-0500

Let "\theta=\theta(t)[°C]" be temperature of a component.

Let the component be generating heat as it is used at rate

"q=q(t)[J\/s]"


Air is blown over the component to cool it.

Temperature of air "\theta_a=\theta_a(t)[°C]"


Temperature of air "\theta_a(t)" and heat generated by component , "q(t)" affect temperature of component "\theta(t)"


The dynamic model for the output "\theta(t)" in response to "q(t)" and "\theta_a(t)" is



"C\>\frac{d\theta}{dt}=q(t)+UA[\theta_a(t)-\theta(t)]"


Where C, U and A are [Constant] parameters.